# On Discounted Infinite-Time Mean Field Games

### v4

@@ 3 Infinite-time mean field games with discounting
In this section, we introduce the infinite-time mean field games with discounting. Let $r>0$ represent the time discount factor and $A\subset\mathbb{R}$ be a convex control space. Define $\mathcal{A}\triangleq L^{2}_{r}(0,\infty,A)$ to be the space of all adimissible controls, and $b,f:\mathbb{R}\times{\cal P}_{2}\times A\rightarrow\mathbb{R}$ are two measurable functions.
@@
where $\beta\in\mathcal{A}$ is the strategy remains to be determined. The representative player seeks to minimize the cost
@@
Assuming $\beta^{x}\in\mathcal{A}$ minimizes the cost function, the state process of the representative player is $X^{x,\beta^{x}}_{t}$. Since the representative agent can characterize the strategies of all players in the population, we assert that the following fundamental relationship must hold:
@@
To solve this mean field game problem, we work under the following assumptions:
@@
These assumptions jointly ensure that: (i) the state process remains confined within the $L^{2}_{r}$ space, and (ii) the cost functional maintains integrability.
Then we partition all agents into two components: (i) the representative player, who dynamically optimizes their strategy, and (ii) the social equilibrium (or mean field), characterizing the macroscopic state shared by the population. For the social equilibrium, its state is governed by the following SDE:
@@
where $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$. We note that, by assumption 3.1, this SDE has a unique strong solution in $L^{2}_{r}$, see ([1], Proposition 2.2) for more details.
@@
We say a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ is a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
@@
When this Nash equilibrium $\alpha^{*}$ exists, we have
@@
Note that our definition of Nash equilibrium differs from that in [1] and earlier finite-time mean field games. In [1], the mean field game is formulated as a fixed-point problem. The population is assumed to be homogeneous, meaning all agents are identical and thus represented by a single representative player. For a given measure flow $(\mu_{t})_{t\geq 0}$ , the representative player wants to minimize
$J^{\mu}(\alpha)\triangleq\mathbb{E}\left[\int_{0}^{\infty}e^{-rt}f(t,X_{t},\mu_{t},\alpha_{t})\mathrm{d}t\right],$ (3.13)
under the constraint
Then we require that the law of $X_{t}$ coincides with $\mu_{t}$, which means we need to find a fixed point.
In this paper, we separate a representative player from the population, who only needs to consider their own optimization problem starting from state $x$. The representative player’s state evolution depends on both their own state $X^{x,\beta}$ and the overall population distribution $\mathcal{L}_{X^{\xi,\alpha}}$. Here we would like to emphasize that since there are a large number of players, any change of a representative player doesn’t impact the measure flow $\mathcal{L}_{X^{\xi,\alpha}}$. Under the existence assumption of the Nash equilibrium $\alpha^{*}$ specified in Definition 3.2, the stochastic dynamics of both the population and representative player are characterized by the following stochastic differential equations:
@@ 4 Connection with infinite-time McKean-Vlasov FBSDEs
In this section, we employ the maximum principle to solve the optimization problem for the representative player and then use the infinite-time McKean-Vlasov FBSDE to construct the optimal strategy for the representative player such that the controls of the representative player and the social equilibrium share the same feedback form. Our derivation is based on the following key assumptions on $b,f$:
@@
(i) $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$. Where $b_{0},f_{0}$ are measurable functions on $\mathbb{R}\times{\cal P}_{2}$, $b_{1},f_{1}$ are measurable functions on $\mathbb{R}\times A$.
(ii) $b,f$ are differentiable with respect to $(x,a)$ and $\partial_{a}b,\partial_{a}f$ are Lipschitz continuous in $(x,a)$.
(iii) $H_{0}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)$ is convex with respect to $(x,a)$. $\min_{a\in A}H_{0}(x,\mu,a,y)$ has a unique minimizer $\hat{\alpha}(x,y)$ which is Lipschitz continuous in $(x,y)$.
@@ 4.1 Pontryagin’s maximum principle
Assuming the state of the social equilibrium $X_{t}^{\xi}$ is given, we consider the optimization problem for the representative player, whose state is given by
@@
Let us formally derive the maximum principle for the infinite-time control problem. Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ . Let
@@
be the variation process. Then it can be shown that $R$ satisfies
@@
and introduce the adjoint process which is determined by an infinite-time BSDE
@@
Applying Itô’s formula to the process $(e^{-rt}R_{t}Y_{t}^{x,\beta})$, by simple computation we can deduce that
@@
Recalling our convexity assumptions on $b,f$ in Assumption (4.1), we know that the representative player’s optimal control $\beta$ takes a feedback form, that is
@@
Now we consider the following two McKean–Vlasov FBSDEs:
@@
Here (4.12) and (4.13) denote the state processes of social equilibrium and the representative player, respectively. Observe that their admissible controls both take the identical feedback form $\hat{\alpha}(x,y)$. We shall prove that when social equilibrium employ this feedback control, the representative player’s loss function is minimized if they use the same feedback form—thereby constituting a Nash equilibrium.
@@
Let $(b,f)$ be differentiable in $(x,a)$, $\mathcal{H}$ be convex in $(x,a)$. Suppose that $\hat{\alpha}$ is Lipschitz continuous and that both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$. If we denote $\hat{\alpha}(X_{t}^{x},Y_{t}^{x})$ as $\alpha^{*}_{t}$ which is an admissible control in $\mathcal{A}$, then we have
@@
Since $X_{t}^{x},X_{t}^{x,\beta},Y_{t}^{x}$ are all belong to $L^{2}_{r}$, we can find a sequence of $T_{i}\rightarrow\infty$ such that
@@
for all admissible control $\beta$. Thus we complete the proof.
@@ 4.2 Solvability of mean field game FBSDEs
In this subsection, we will find sufficient conditions for the existence and uniqueness of solutions to (4.12) and (4.13). Considering the linear case, we assume $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
We require a technical lemma about the Lipschitz and convex property of the minimizer $\hat{\alpha}$, the detailed proof of which can be found in ([1], Lemma 3.1).
@@
Suppose $f_{1}$ is once continuously differentiable in $(x,a)$, $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$. And $f_{1}$ is $\eta$-convex in $a$, which means
@@
Then it holds that
@@
Furthermore, if $A=\mathbb{R}$ and $\partial_{a}f$ is $\zeta$-Lipchitz in $a$, it follows that
@@
Let $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$. Under the following conditions, Assumption 3.1 and 4.1 are satisfied, and both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
@@
(ii) $f_{0}(x,\mu)$ is convex in $x$ and there exists a positive constant $b_{4}>0$, such that
$|\partial_{x}f_{0}(x,\mu)-\partial_{x}f_{0}(x^{\prime},\mu^{\prime})|<b_{4}\left(|x-x^{\prime}|+W_{2}(\mu,\mu^{\prime})\right).$ (4.25)
(iii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds
@@
(iv) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipchitz continuous in $(x,a)$.
@@
If we set $f_{0}(x,\mu)=b_{4}x\bar{\mu},f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C,\ell=0$. Then the requirement in (4.27) becomes
@@
Proof.  It’s clear that Assumption 3.1 is satisfied. And by Lemma 4.3, $\hat{\alpha}$ is Lipchitz. In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+f_{0}(x,\mu)+f_{1}(x,a)$ is convex in $(x,a)$ according to assumption on $f_{0},f_{1}$, thus Assumption 4.1 is satisfied. To prove that the BSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$, the only condition that remains to be verified is (2.5).
We consider the FBSDE (4.12), and set
@@
Take four arbitrary square integrable random variables $X,Y,X^{\prime},Y^{\prime}$. Define $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}})$ . We have
@@ 5 Master equation representation
Assume the master equation (1.7) has a classical solution $U(x,\mu)\in{\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$ with $H$ defined above, and $\hat{b}(x,\mu)=b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$. For any $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$, suppose the following SDE admits a unique solution in $L^{2}_{r}$:
@@
Then we take $\mathcal{Y}_{t}^{\xi}=\partial_{x}U(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}})$. By differentiating both sides of the master equation (1.7) with respect to $x$ and applying Itô’s formula, we obtain
@@
Assume the master equation (1.7) admits a classical solution $U(x,\mu)\in{\cal C}^{2,1}(\mathbb{R}\times{\cal P}_{2})$ which is of at most quadratic growth, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$:
@@
Then for every initial distribution $\mu_{0}$, $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium, and $U(x,\mu_{0})$ is exactly the value function given $\alpha^{*}$.
Proof.  Given $\alpha^{*}$ and initial $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu_{0})$, the states of social equilibrium is governed by the following SDE:
@@
for every feasible control $\beta$, and the equality holds when $\beta_{t}=\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t}))$. This shows $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is an Nash equilibrium. And we have
$U(x,\mu_{0})=\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t})))\mathrm{d}t,$ (5.13)
this completes the proofs of our desired results.
At the end of this section, we provide an example of a solvable elliptic master equation. Set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a,f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, where $b_{1},b_{2},b_{3},b_{4},A,C$ are constants satisfying all conditions in Theorem 4.4 and Remark 4.5. Then we have $\hat{\alpha}(x,y)=-\frac{b_{3}y}{2C}$ and
@@
Theorem 5.2
The master equation (5.15) admits a classical solution $U(x,\mu)$ in ${\cal C}^{3,1}(\mathbb{R}\times{\cal P}_{2})$.
Proof.  We assume that the solution takes the form
@@
where $a_{1},a_{2},a_{3},a_{4}$ are constants remain to be determined. Then we have
@@
Then, we can solve for $a_{1},a_{2},a_{3},a_{4}$.
Remark 5.3
We have
$\hat{b}(x,\mu)\triangleq b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))=(b_{1}-\frac{a_{1}b_{3}^{2}}{C})x+(b_{2}-\frac{a_{2}b_{3}^{2}}{2C})\bar{\mu}.$ (5.20)
Since the SDE (5.3) now becomes linear, if we further require $a_{1}>0,a_{2}>0$, it’s not hard to see that the SDE admits a unique solution in $L_{r}^{2}$.

### v5

@@ 3 Infinite-time mean field games with discounting
In this section, we introduce infinite-time mean field games with discounting. Let $r>0$ represent the time discount factor and $A\subset\mathbb{R}$ be a convex control space. Define $\mathcal{A}\triangleq L^{2}_{r}(0,\infty,A)$ to be the space of all admissible controls, and $b,f:\mathbb{R}\times{\cal P}_{2}\times A\rightarrow\mathbb{R}$ are two measurable functions.
@@
where $\beta\in\mathcal{A}$ is the strategy that remains to be determined. The representative player seeks to minimize the cost
@@
Assuming that $\beta^{x}\in\mathcal{A}$ minimizes the cost function, the state process of the representative player is $X^{x,\beta^{x}}_{t}$. Since the representative player can characterize the strategies of all players with the same initial state in the population, we assert that the following fundamental relationship must hold:
@@
To solve this mean field game problem, we work under the following assumptions.
@@
These assumptions jointly ensure that (i) the state process remains confined within the $L^{2}_{r}$ space and (ii) the cost functional maintains integrability.
Then, we partition all agents into two components: (i) the representative player, who dynamically optimizes their strategy, and (ii) the social equilibrium (or mean field), characterizing the macroscopic state shared by the population. The state of the social equilibrium is governed by the following stochastic differential equation (SDE):
@@
where $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $\alpha\in\mathcal{A}$. We note that by assumption 3.1, this SDE has a unique strong solution in $L^{2}_{r}$; see [1] (Proposition 2.2) for more details.
@@
We say that a Lipschitz function ${\alpha}^{*}(x,\mu):\mathbb{R}\times{\cal P}_{2}\rightarrow\mathbb{R}$ constitutes a discounted infinite-time mean field Nash equilibrium for a given initial distribution $\mu_{0}$ if for any initial state $\xi_{0}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with distribution $\mu_{0}$, the closed-loop controls $\alpha_{s}^{\xi_{0}}={\alpha}^{*}(X_{s}^{\xi_{0},\alpha^{\xi_{0}}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$$\alpha_{s}^{x}={\alpha}^{*}(X_{s}^{x,\alpha^{x}},\mathcal{L}_{X_{s}^{\xi_{0},\alpha^{\xi_{0}}}})$ satisfy
@@
When the Nash equilibrium $\alpha^{*}$ exists, we have
@@
In our framework, we separate a representative player from the population, who only needs to consider their optimization problem starting from state $x$. This model decouples the micro-level agent from the macro-level population distribution, enabling an interconnected analysis of their evolution. The equilibrium is characterized by two consistency conditions.
  • Macro-consistency: The aggregate distribution generated by all players adopting this strategy must coincide with the posited social equilibrium.
The representative player’s state evolution depends on both their state $X^{x,\beta}$ and the overall population distribution $\mathcal{L}_{X^{\xi,\alpha}}$. Here, we would like to emphasize that since the number of players is large, any change by a representative player does not impact the measure flow $\mathcal{L}_{X^{\xi,\alpha}}$. Under the existence assumption of the Nash equilibrium $\alpha^{*}$ specified in Definition 3.2, the stochastic dynamics of both the population and representative player are characterized by the following SDE:
@@ 4 Connection with infinite-time McKean-Vlasov FBSDEs
In this section, we employ the maximum principle to solve the optimization problem for the representative player and then use infinite-time McKean-Vlasov FBSDEs to construct the optimal strategy for the representative player such that the controls of the representative player and the social equilibrium share the same feedback form. Our derivation is based on the following key assumptions on $b,f$:
@@
(i) $b(x,\mu,a)=b_{0}(x,\mu)+b_{1}(x,a)$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$ where $b_{0},f_{0}$ are measurable functions on $\mathbb{R}\times{\cal P}_{2}$ and $b_{1},f_{1}$ are measurable functions on $\mathbb{R}\times A$.
(ii) $b,f$ are differentiable with respect to $(x,a)$, and $\partial_{a}b,\partial_{a}f$ are Lipschitz continuous in $(x,a)$.
(iii) $H_{0}(x,\mu,a,y)\triangleq b(x,\mu,a)\cdot y+f(x,\mu,a)$ is convex with respect to $(x,a)$. $\min_{a\in A}H_{0}(x,\mu,a,y)$ has a unique minimizer $\hat{\alpha}(x,y)$ that is Lipschitz continuous in $(x,y)$.
@@ 4.1 Pontryagin’s maximum principle
Assuming that the state of the social equilibrium $X_{t}^{\xi}$ is given, we consider the optimization problem for the representative player, whose state is given by
@@
Let us formally derive the maximum principle for the infinite-time control problem. Suppose $\beta$ is an optimal control, choose another admissible control $\gamma$, and denote by $X^{x,\beta+\epsilon\gamma}$ the state trajectory corresponding to the control $\beta+\epsilon\gamma$ . Let
@@
be the variation process. Then, it can be shown that $R$ satisfies
@@
and introduce the adjoint process, which is determined by an infinite-time BSDE:
@@
Applying Itô’s formula to the process $(e^{-rt}R_{t}Y_{t}^{x,\beta})$ and by simple computation, we can deduce that
@@
Recalling our convexity assumptions on $b,f$ in Assumption 4.1, we know that the representative player’s optimal control $\beta$ takes a feedback form, that is,
@@
Now, we consider the following two McKean-Vlasov FBSDEs:
@@
Here, (4.12) and (4.13) denote the state processes of the social equilibrium and the representative player, respectively. Observe that their admissible controls both take the identical feedback form $\hat{\alpha}(x,y)$. We shall prove that when the social equilibrium employs this feedback control, the representative player’s loss function is minimized if they use the same feedback form, thereby constituting a Nash equilibrium.
@@
Let $(b,f)$ be differentiable in $(x,a)$ and $\mathcal{H}$ be convex in $(x,a)$. Suppose that $\hat{\alpha}$ is Lipschitz continuous and that both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$. If we denote $\hat{\alpha}(X_{t}^{x},Y_{t}^{x})$ as $\alpha^{*}_{t}$, which is an admissible control in $\mathcal{A}$, then we have
@@
Since $X_{t}^{x},X_{t}^{x,\beta},Y_{t}^{x}$ all belong to $L^{2}_{r}$, we can find a sequence of $T_{i}\rightarrow\infty$ such that
@@
for all admissible controls $\beta$. Thus, we complete the proof.
@@ 4.2 Solvability of mean field game FBSDEs
In this subsection, we will find sufficient conditions for the existence and uniqueness of solutions to (4.12) and (4.13). Considering the linear case, we assume that $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$, where $\bar{\mu}$ is the mean value of the probability measure $\mu$ and $b_{1},b_{2},b_{3},b_{4}$ are constants.
We require a technical lemma about the Lipschitz and convex properties of the minimizer $\hat{\alpha}$, the detailed proof of which can be found in [1] (Lemma 3.1).
@@
Suppose $f_{1}$ is once continuously differentiable in $(x,a)$, $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $f_{1}$ is $\eta$-convex in $a$, which means that
@@
Then, it holds that
@@
Furthermore, if $A=\mathbb{R}$ and $\partial_{a}f$ is $\zeta$-Lipschitz in $a$, it follows that
@@
Let $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=f_{0}(x,\mu)+f_{1}(x,a)$. Under the following conditions, Assumptions 3.1 and 4.1 are satisfied and both (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
@@
(ii) $f_{0}(x,\mu)$ is convex in $x$, and there exists a positive constant $b_{4}>0$ such that
$|\partial_{x}f_{0}(x,\mu)-\partial_{x}f_{0}(x^{\prime},\mu^{\prime})|<b_{4}\left(|x-x^{\prime}|+{\cal W}_{2}(\mu,\mu^{\prime})\right).$ (4.25)
(iii) There exist some positive constants $\eta,\iota$ such that the following convexity condition holds:
@@
(iv) There exist some positive constants $\zeta,\ell$ such that $\partial_{a}f_{1}$ is $\ell$-Lipschitz in $x$ and $\zeta$-Lipschitz in $a$. $\partial_{x}f_{1}$ is Lipschitz continuous in $(x,a)$.
@@
If we set $f_{0}(x,\mu)=b_{4}x\bar{\mu},f_{1}(x,a)=Ax^{2}+Ca^{2},A>0,C>0$, we have $\iota=A,\eta=C,\zeta=2C,\ell=0$. Then, the requirement in (4.27) becomes
@@
Proof.  It is clear that Assumption 3.1 is satisfied. And by Lemma 4.3, $\hat{\alpha}$ is Lipschitz. In addition, $\mathcal{H}(x,\mu,a,y)=(b_{1}x+b_{2}\bar{\mu}+b_{3}a)\cdot y+f_{0}(x,\mu)+f_{1}(x,a)$ is convex in $(x,a)$ according to the assumption on $f_{0},f_{1}$; thus, Assumption 4.1 is satisfied. To prove that the FBSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$, the only condition that remains to be verified is (2.5).
We consider the FBSDE (4.12) and set
@@
Take four arbitrary square-integrable random variables $X,Y,X^{\prime},Y^{\prime}$. Define $\hat{X}=X-X^{\prime},\hat{Y}=Y-Y^{\prime}$ and $U=(X,Y,\mathcal{L}_{X}),U^{\prime}=(X^{\prime},Y^{\prime},\mathcal{L}_{X^{\prime}})$ . We have
@@
By Lemma 2.2, we know both FBSDEs (4.12) and (4.13) admit unique strong solutions in $L^{2}_{r}$.
@@ 5 Master equation representation
Assume that the master equation (1.8) has a classical solution $U(x,\mu)\in C^{2,1}(\mathbb{R}\times{\cal P}_{2})$ with $H$ defined above such that $\partial_{x}U(x,\mu)\in C^{2,1}(\mathbb{R}\times{\cal P}_{2})$ satisfying PDE (1.12) and $\hat{b}(x,\mu)=b(x,\mu,\hat{\alpha}(x,\partial_{x}U(x,\mu)))$ is Lipschitz continuous in $(x,\mu)$. For any $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$, suppose the following SDE admits a unique solution in $L^{2}_{r}$:
@@
Then, we take $\mathcal{Y}_{t}^{\xi}=\partial_{x}U(\mathcal{X}_{t}^{\xi},\mathcal{L}_{\mathcal{X}_{t}^{\xi}})$. By differentiating both sides of the master equation (1.8) with respect to $x$ and applying Itô’s formula, we obtain
@@
Assume that the master equation (1.8) admits a classical solution $U(x,\mu)\in C^{2,1}(\mathbb{R}\times{\cal P}_{2})$, which is of at most quadratic growth, and the following SDEs admit unique solutions in $L^{2}_{r}$ for each $x\in\mathbb{R}$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu)$
@@
Then, for every initial distribution $\mu_{0}$, $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium, and $U(x,\mu_{0})$ is exactly the value function given $\alpha^{*}$.
Proof.  Given $\alpha^{*}$ and initial $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mu_{0})$, the state of social equilibrium is governed by the following SDE:
@@
for every feasible control $\beta$, and the equality holds when $\beta_{t}=\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t}))$. This shows that $\alpha^{*}(x,\mu)=\hat{\alpha}(x,\partial_{x}U(x,\mu))$ is a Nash equilibrium. In addition, we have
$U(x,\mu_{0})=\mathbb{E}\int_{0}^{+\infty}\mathrm{e}^{-rt}f(X^{x}_{t},\rho_{t},\hat{\alpha}(X_{t}^{x},\partial_{x}U(X_{t}^{x},\rho_{t})))\mathrm{d}t.$ (5.13)
This completes the proof of our desired results.
At the end of this section, we provide an example of a solvable elliptic master equation and demonstrate that its solutions are generally nonunique. Set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a,f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, where $b_{1},b_{2},b_{3},b_{4},A,C$ are constants satisfying all conditions in Theorem 4.4 and Remark 4.5. Then, we have $\hat{\alpha}(x,y)=-\frac{b_{3}y}{2C}$ and
@@
We assume that the solution takes the form
@@
where $a_{1},a_{2},a_{3},a_{4}$ are constants that remain to be determined. Then we have
@@
Note that $a_{1},a_{2},a_{3}$, and $a_{4}$ can be solved sequentially, and the equations for $a_{1}$ and $a_{2}$ are quadratic. Since $A>0$ and $\frac{b_{3}^{2}}{C}<0$, the equation for $a_{1}$ must have one positive and one negative root. If the coefficients are sufficiently well-behaved, for instance, $2b_{2}a_{1}+b_{4}>0$, then the equation for $a_{2}$ will also have two distinct real roots. Therefore, the system of equations (5.19) can have at most four solutions. It must be emphasized that there exists at most one set of solutions for which the solution of SDE (5.3) lies in $L_{r}^{2}$, as the $L_{r}^{2}$ solution of FBSDE (4.12) is unique. In a companion article, we establish the uniqueness of solutions to the master equation under certain monotonicity and growth assumptions.
Example 5.2
We consider a case that $r=2,b_{1}=b_{2}=b_{4}=0,b_{3}=2,A=2,C=1$, then equation (5.15) becomes
$2U(x,\mu)=$ $2x^{2}-\left(\partial_{x}U(x,\mu)\right)^{2}+\frac{1}{2}\partial_{xx}U(x,\mu)$ (5.20)
$+\tilde{\mathbb{E}}\left[\frac{1}{2}\partial_{\tilde{x}}\partial_{\mu}U(x,\mu,\tilde{\xi})-2\partial_{\mu}U(x,\mu,\tilde{\xi})\cdot\partial_{x}U(\tilde{\xi},\mu)\right]$
Solving the corresponding Equation (5.19) yields the following solutions:
$U_{1}(x,\mu)=\frac{1}{2}x^{2}+\frac{1}{4},$ (5.21)
$U_{2}(x,\mu)=\frac{1}{2}x^{2}-3x\bar{\mu}+\frac{3}{2}(\bar{\mu})^{2}+\frac{1}{4},$
$U_{3}(x,\mu)=-x^{2}-\frac{1}{2},$
$U_{4}(x,\mu)=-x^{2}+3x\bar{\mu}-\frac{3}{2}(\bar{\mu})^{2}-\frac{1}{2}.$
However, to ensure that the solution of Equation (5.3) , now written as
$\mathrm{d}X_{t}=-4\partial_{x}U(X_{t},{\cal L}_{X_{t}})\mathrm{d}t+\mathrm{d}B_{t},\quad X_{0}=\xi,$ (5.22)
belong to $L_{2}^{2}$, only the solution $U_{1}(x,\mu)=\frac{1}{2}x^{2}+\frac{1}{4}$ satisfies this condition.